OpenStudy (anonymous):

Which of the following shows the step necessary to solve the equation x plus two and one-half equals five and five-sixths and its correct solution? Subtract two and one-half from both sides; three and one-third Add two and one-half to both sides; eight and one-third Subtract five and five-sixths from both sides; three and one-third Add five and five-sixths to both sides; eight and one-third

OpenStudy (anonymous):

@nikato @dg2 @GhostKillerPlay

OpenStudy (anonymous):

wait:)

OpenStudy (anonymous):

x +2 1/2 = 5 5/6 x +2 1/2 -2 1/2 = 5 5/6 -2 1/2 x = 5 5/6 - 2 3/6 x = 3 2/6 or 3 1/3

OpenStudy (anonymous):

@dg2 what's the answer?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Subtract two and one-half from both sides; three and one-third

OpenStudy (anonymous):

x +2 1/2 = 5 5/6

OpenStudy (anonymous):

Subtract two and one-half from both sides; three and one-third

OpenStudy (anonymous):

we get 3 1/3